Knowledge technologies include computer interfaces and networked systems that enable organizational knowledge exchanges among human actors. Organizational stimuli include formal incentives, normative values, and competencebased trust to motivate human actors to transfer knowledge intra- and interorganizationally. Knowledge exchange processes embody intra- and inter-organizational knowledge creation, transfer, and protection activities.

This seminar will examine one aspect of knowledge ecosystems in-depth, specifically that of structure and ”top-down” vs. ”bottom-up” knowledge management strategies by extending March’s (1991) organizational science model of exploration and exploitation to demonstrate how these elements, given environmental turbulence, influence aggregate individual knowledge levels within an organization. While other elements of my dissertation examine organizational stimuli and knowledge exchange processes, this seminar will focus specifically on organizational design and strategy elements. Moreover, while this seminar will be given to a predominantly computer science audience the audience is given fair warning that this seminar focuses on information systems, to include the human factors associated with system use.